OpenStudy (anonymous):

critical points for f(t)=8t^3-t^2

OpenStudy (anonymous):

To find critical points you take the derivative and set it equal to zero: f'(t) = 24t^2-2t 0 =t(24t - 2) t = 0 or 1/12

OpenStudy (turingtest):

\[ f(t)=8t^3-t^2\]\[ f'(t)=24t^2-2t=t(24t-2)=0\]\[t=\left\{ 0,1/12 \right\}\] continuing for the second derivative...

OpenStudy (turingtest):

\[f''(t)=48t-2=0\]\[t=\left\{ 1/24 \right\}\]